Eggz, Inc., is considering the purchase of new equipment that will allow the company to collect loose hen feathers for sale. The equipment will cost $425,000 and will be eligible for 100 percent bonus depreciation. The equipment can be sold for $45,000 at the end of the project in five years. Sales would be $275,000 per year, with annual fixed costs of $47,000 and variable costs equal to 35 percent of sales. The project would require an investment of $25,000 in NWC that would be returned at the end of the project. The tax rate is 22 percent, and the required return is 9 percent. What is the project’s NPV?

PLEASE SHOW EXCEL FUNCTIONS
Answer #1

Operating cash flow (OCF) each year = income after tax + depreciation

In year 5, the entire NWC is recovered.

profit on sale of equipment at end of year 5 =  sale price - book value

book value = original cost - accumulated depreciation

accumulated depreciation = zero, as the equipment is fully depreciated in the 1st year itself.

after-tax salvage value = salvage value - tax on profit on sale of equipment   

NPV is calculated using NPV function in Excel

NPV is $99,561

$425,000 $25,000 $275,000 $47,000 $96,250 $275,000 $47,000 $96,250 $275,000 $47,000 $96,250 $275,000 $47,000 $96,250 2 Initia

В In 4 2 Initial Investment 3 Cost of new equipment 4 Increase in NWC 425000 25000 275000 47000 =D7*35% 275000 47000 =E7*35%